Friday 03/23/12

10:45 A – 12:30 A

  • O/N confocal 10 um is a little too much buffer between first and last, several embryos didn’t make it all the way through.  Try again tonight.
  • Set up new O/N confocal run 6 um difference in initial frame, 60 frames to improve chance of getting everything.
  • Psc, Pc, Su(Z), Pho with K4 and with K27 in EtOH. Then embedding into resin.
  • Psc, Pc, Su(Z), Pho with HP1 rinsing out HP1 secondary in PBS.  Then add WGA-488 1 hr.  Then rinse out WGA 45 min 4 washes.  post fix.
  • Bcd, en-locus BH treat, and en-locus regular all in PBT after rinsing out O/N H3-750.  Also WGA-488 treated 45 min 4 washes and post fix 1 hr.
  • LB amp plates from cloning all lawns.  Streak some cells to open spots on plate.  Order beads for coating.
  • Make up new LB amp and pick colonies from cloning.
Confocal testing
  • Bcd hb-y(i) looks beautiful.
  • Does have weak 647 nuclear dots (are these the ‘PcG bodies’?)
  • Bmi antibody is mouse, works well in lightly fixed samples
  • co-stained BMI with rb a-647, still has nuclear bodies.  Might be this particular batch of antibody?
  • Test sample of Psc K27 647 channel looks pretty weak, membranes don’t look good — degradation too short a fix?
  • MeOH stored PcG proteins stains look strong, but no nuclear bodies. These may have been an artifact of one of the rb antibodies in short fixed embryos.
  • original eng probe set does work: en-dnp, inv-dnp, E(Pc)-dig, tou-bio.  dots not especially bright, bio channel and red channel backgroundry.
  • Early sodium borohydride does not seem to have made obvious staining difference.


  • Laser power issues: 750 max out at 80 mW, 647 max out at less than 50 mW.
  • Graham fixes lasers, 750 now 90 mW, 647 now 100 mW.
  • Imaging Psc and K27me3 in anterior cells at germ-band extension and germband rectraction phases.


  • Send review papers to XZ
  • Send ACS paperwork to XZ



This entry was posted in Summaries and tagged , , . Bookmark the permalink.